OpenStudy (anonymous):

Can someone help me solve these and explain the steps to me 1. 2/(x+3) - 1/x = -6/x(x+3) 2. 1/2 + x/6 = 18/x

OpenStudy (anonymous):

@radar @satellite73

OpenStudy (anonymous):

Let's do problem 1 first. We need to get rid of the fractions. Do you know how we would do that?

OpenStudy (anonymous):

No :( don't we need to find the LCD? @nelsonjedi

OpenStudy (anonymous):

Yes....so waht is the LCD?

OpenStudy (anonymous):

x +3?

OpenStudy (anonymous):

Will x+3 go into x?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

The LCD is the lowest factor of the numbers. Thus x(x+3) is the LCD.

OpenStudy (anonymous):

ooh okay so now that we determine x(x+3) as the LCD what do we do? I got this is this right? 2x/x(x+3) - 1(x+3)/x(x+3) = -6/x(x+3)

OpenStudy (anonymous):

You multiply each number by the LCD. So let's do 2 / (x+3) if I multiply it by x(x+3) what do I get?

OpenStudy (anonymous):

why would you multiply by the whole LCD, don't you just want to get the denominator equal to the LCD?

OpenStudy (anonymous):

No we want to get rid of the denominator when solving fractions. So what is 2/(x+3) * x (x+3) =?

OpenStudy (anonymous):

2x(x+3)/x^2 + 6x + 9

OpenStudy (anonymous):

\[\frac{ 2 }{ (x+3) } X \frac{ x(x+3) }{ 1 }\]

OpenStudy (anonymous):

This is how the equation would look. so we would end up with 2 * x * (x+3) / (x+3)...Correct?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

The (x+3) cancels each other out so we end up with what?

OpenStudy (anonymous):

2x!

OpenStudy (anonymous):

Bingo. Now can you do the second one for 1/x?

OpenStudy (anonymous):

no 2x^2

OpenStudy (anonymous):

ooh nvm lol wouldn't you get just x?

OpenStudy (anonymous):

sorry for 1/x wouldn't you get x?

OpenStudy (anonymous):

Multiply the top by the LCD

OpenStudy (anonymous):

Ohh you would get (x+3)

OpenStudy (anonymous):

Correct. So what would it be for 6/(x)(x+3) ?

OpenStudy (anonymous):

you would get 6 because everything would cancel out right?

OpenStudy (anonymous):

I don't get what to do with 2x - x+3 = 6 ?

OpenStudy (anonymous):

I don't get what to do with 2x - x+3 = 6 ?

OpenStudy (radar):

Just maybe there are no solution to these!

OpenStudy (anonymous):

@radar I know there is no answer to the first one but the second one of -12, 9 I got the answers from thee back of my book, but I just want to know how to solve them :/

OpenStudy (radar):